I am in the sixth grade, I need help with this equation. 24x - 10 = 18x - 4. thanks, Skyla

Found 2 solutions by solver91311, Edwin McCravy:
Answer by solver91311(24713)



Add -18x to both sides

Add 10 to both sides

Divide by 6

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Edwin McCravy(20055)

 24x - 10 = 18x -  4

Get rid of the - 10 on the left with + 10 on both sides:

 24x - 10 = 18x -  4
     + 10       + 10
 24x      = 18x +  6

Get rid of the 18x on the right with -10x on both sides:

 24x - 10 = 18x -  4
     + 10       + 10
 24x      = 18x +  6
-18x       -18x    
  6x      =        6

Get rid of the 6 on the left by dividing both sides by 6

 24x - 10 = 18x - 4
     + 10       + 10
 24x      = 18x + 6
-18x       -18x    
  6x      =       6
  6               6

Divide the 6's

 24x - 10 = 18x - 4
     + 10       + 10
 24x      = 18x + 6
-18x       -18x    
  6x      =       6
  6               6
   x      =       1 


Edwin